OpenStudy (anonymous):

The length of a rectangle is three feet shorter than twice it's width. If the area of the rectangle is 170ft^(2) , What is the length of the rectangle? Thx.

OpenStudy (ash2326):

Let the length of rectangle be \(l\) and width be \(w\) , it's given that length of rectangle is 3 feet shorter than twice the width so \[l=2w-3\] Area is given as 170 ft^2 so \[l w= 170\] we have l= 2w-3, so \[(2w-3)(w)=170\] \[2w^2-3w=170\] we have now \[2w^2-3w-170=0\] now we know that solution of a quadratic equation \[ax^2+bx+c=0\] is given by \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] now we have \[2w^2-3w-170=0\] here a=2 b=-3 and c=-170 so we get \[w=\frac{-(-3)\pm \sqrt{(-3)^2-4\times 2 \times -170}}{2\times 2}\] we get \[w=\frac{3\pm \sqrt{9+1360}}{4}\] we have now [\[w=\frac{3\pm \sqrt{1369}}{4}\] we know that \(\sqrt{1369}=37\) so we get \[w=\frac{3\pm 37}{4}\] either \[w=\frac{3+ 37}{4}or \frac{3- 37}{4}\] since width can't be negative, we'llhave \[w= \frac{3+37}{4}\] we get \[w=10\] now we know \[lw=170\] w=10 so \[l\times 10=170=>l=17\] so we have l=17 feet and w=10 feet

OpenStudy (ash2326):

@leidyr check this out, do you understand this?

OpenStudy (anonymous):

Yes, well explained it. Thanks so much..